OpenStudy (anonymous):

find the slope between (1,6) and (-8, 12)

OpenStudy (anonymous):

slope is\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

that is the formula

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i already know that just confused on how to do it lol

OpenStudy (anonymous):

\[\frac{ 12-6 }{ -8-1 }\]

OpenStudy (anonymous):

now just subtract

OpenStudy (anonymous):

6/-7

OpenStudy (anonymous):

6 is right

OpenStudy (anonymous):

oh 9

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

it would be -9

OpenStudy (anonymous):

yes I forgot the negative sign you must be a whiz at math thanks

OpenStudy (anonymous):

haha thanks and no problem :)

OpenStudy (anonymous):

y=5x-7 m= b=

OpenStudy (anonymous):

Im not sure on this one sorry

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

I would post that question up

OpenStudy (anonymous):

m= and y=

OpenStudy (mathstudent55):

Compare y = 5x - 7 with y = mx + b Can you see what m and b are?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

x+y =0

OpenStudy (anonymous):

would it be 1

OpenStudy (mathstudent55):

Are you looking for m and b for x + y = 0?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Then remember how we looked at y = 5x - 7, and we directly compared to y = mx + b, so we can conclude m = 5 and b = -7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now you have x + y = 0. If you solve for y, and express it as y = ... You'l be able to again get m and b. First solve x + y = 0 for y. What do you get?

OpenStudy (anonymous):

1

OpenStudy (mathstudent55):

No. I don't want m yet. Do one step at a time. If you have equation x + y = 0, solve it for y. You need to subtract x from both sides.

OpenStudy (anonymous):

so O is the answer Zero or undefined

OpenStudy (mathstudent55):

I give up. I've asked you to solve the equation x + y = 0 for y, a few times. I told you you need to subtract x from both sides. If you don't want to follow what I ask, I can't help you.

OpenStudy (anonymous):

Im asking you to explain what to do because Im not sure how to do the problem or write it x-x + y = 0-x

OpenStudy (mathstudent55):

Ok, I'll explain it now that I see you are trying to solve for y.

OpenStudy (mathstudent55):